# Continuum Attention for Neural Operators

### v1

@@ 5 Universal Approximation by Transformer Neural Operators
The existence of $W_{1},R_{1},R_{2},P_{1},P_{2},b_{1},b_{R},b_{R}^{\prime},b_{P},b_{P}^{\prime}$ so that $\mathsf{G}(\cdot\,;\theta)$ satisfies (35) then follows from Lanthaler et al. (2023, Theorem 2.1), which also involves the application of the universality result for two-layer neural networks of Pinkus (1999, Theorem 4.1).
The analysis in Lanthaler et al. (2023) allows the derivation of an analogous universal approximation theorem for functions belonging to Sobolev spaces. This generalization is the content of the next theorem. The proof follows easily by applying the same argument as in the proof of Theorem 22 and the result of Lanthaler et al. (2023, Theorem 2.2).
@@ 6 Numerical Experiments
In this section we illustrate, through numerical experiments, the capability of the transformer neural operator architectures described in Section 4. Throughout, we consider the supervised learning problem described by the data model in (3) and take the viewpoint of surrogate modeling to construct approximations of the operators $\mathsf{G}^{\dagger}$. In Subsection 6.1 we consider problems given by dynamical systems and ordinary differential equations, namely operators acting on functions defined on a one-dimensional time domain. In this context, we explore the use of the transformer neural operator for operator learning problems given by the Lorenz ‘63 dynamical system and a controlled ODE. On the other hand, in Subsection 6.2 we consider problems given by partial differential equations equations, namely operators acting on functions defined on a two-dimensional spatial domain. In this context we explore the use of the transformer neural operator and of the more efficient ViT neural operator and Fourier attention neural operator architectures for operator learning problems given by the Darcy flow and Navier-Stokes equations.
@@
${dz}$ $=\bigl{(}xy-\beta z\bigr{)}{dt}.$
We consider learning the operator $\mathsf{G}^{\dagger}:C\bigl{(}[0,T];\mathbbm{R}^{d_{x}}\bigr{)}\times\mathbbm{R}^{d_{y}}\times\mathbbm{R}^{d_{z}}\to C\bigl{(}[0,T];\mathbbm{R}^{d_{y}}\bigr{)}\times C\bigl{(}[0,T];\mathbbm{R}^{d_{z}}\bigr{)}$ given by
@@
The experimental settings we consider for the continuous time operator learning problem may be viewed from the lens of data assimilation as smoothing problems (Sanz-Alonso et al., 2023). Indeed, the mappings to be learned represent the recovery of an unobserved trajectory $\{z(t)\}_{t\in{[0,T]}}$ conditioned on the observation of a whole trajectory $\{u(t)\}_{t\in{[0,T]}}$.
Figure 7: The panel displays the performance in relative $L^{2}$ errors (in log-scale) of the transformer neural operator when applied to the Lorenz ‘63 and controlled ODE operator learning test sets of different resolutions at inference time (without retraining). All models were parametrized by $d_{\textrm{model}}=128$ and $6$ encoder layers, and were trained with $8000$ trajectories with $T=2$, sampled according to $\Delta t=0.01$.
@@
In other words, the second half of the testing domain is up-sampled by a factor of 2 compared to the training discretization. As seen qualitatively in Figure 8 testing on the non-uniform discretization based on a model trained on the uniform grid is successful. The use of different grids in test and train does leads to a larger error than the one incurred when they match (seen in Figure 7), but the errors remain small. It is our continuum formulation of attention from Section B.1 that enables this direct generalization; it would not be possible with standard methods.
@@
In this subsection we study two different 2D operator learning problems – flow in a porous medium, governed by the Darcy model, and incompressible viscous fluid mechanics governed by the Navier-Stokes equation, in particular, a Kolmogorov flow. We apply the three architectures developed in Sections 2 and 3, concentrating on the patched architectures. In the first Subsection 6.2.1 we report the results obtained by applying the neural operators described to an array of operator learning problems, where the domain of the functions is two-dimensional i.e. $D\subset\mathbbm{R}^{2}$. In Subsection 6.2.2 we describe the settings of the two Darcy flow operator learning problems considered, along with the parametrization and training details for the various architectures. Similarly in Subsection 6.2.3, we describe the Kolmogorov flow operator learning problem and relevant implementation details. Subsections 6.2.2 and 6.2.3 also contain further numerical results illustrating the behaviour of the proposed transformer neural operators, and discretization invariance in particular.
@@ 6.2.1 Results
The results presented concern the two variants of Darcy flow, with lognormal and piecewise constant inputs, and the Kolmogorov flow. The operator learning architectures we deploy include the three transformer-based methods introduced in this paper, and the implementations of the Fourier neural operator from the “NeuralOperator” library (Li et al., 2021; Kovachki et al., 2023).
Our first results compares the transformer neural operator from Subsection 4.2 with FNO on the lognormal Darcy Flow problem in a low $64\times 64$ resolution setting. Table 1 provides a brief summary of the test relative $L^{2}$ errors obtained. It is notable that the transformer architecture obtains performance similar to FNO, despite not using patching, and with an order of magnitude fewer parameters. However since higher resolution renders the architecture from Subsection 4.2 impractical in dimensions $d\geq 2$ the remainder of our experiments use the patched ViT and Fourier attention neural operators from Subsections 4.3 and 4.4 respectively.
@@
Transformer NO $5.98\cdot 10^{5}$ $5.29\cdot 10^{-3}$
FNO $1.11\cdot 10^{7}$ $5.74\cdot 10^{-3}$
@@
We now consider training two patched architectures on the Darcy Flow problem with piecewise constant inputs and the Kolmogorov Flow problem; in both cases we use a $416\times 416$ resolution setting. We compare cost versus accuracy for the different neural operators implemented, defining cost in terms of both number of parameters and number of FLOPS. In Table 2 we present the FLOPS for each method; details of the derivation are provided in Appendix C. We note that the parameter scaling for the transformer neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\bigr{)}$, while the scaling for the FNO and the Fourier attention neural operator is $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\cdot k_{\textrm{max}}\bigr{)}$, where $k_{\textrm{max}}$ is the total number of Fourier modes used. The parameter scaling for the ViT neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}\cdot k_{\textrm{max}}+d_{\textrm{model}}^{2}\bigr{)}$. Further details are provided in Appendix C.
@@
Table 2: Evaluation cost (measured in FLOPS) for the three proposed transformer neural operators compared to the Fourier neural operator (Li et al., 2021). We give further details on the derivation of these scalings in Appendix C.
Figures 9 and 10 display the cost-accuracy trade-off. Table 3 provides the test relative $L^{2}$ errors values, depicted in the cost-accuracy analysis figures, obtained with the models with the highest parameter count. Figure 9 demonstrates the parameter-efficiency of the ViT neural operator in comparison with both the FNO and Fourier attention neural operator. Once cost is defined in terms of FLOPS it is apparent that both patched transformer architectures outperform the FNO; furthermore the FNO exhibits accuracy-saturation, caused by the finite data set, whereas the patched architectures are less affected, in the ranges we test here. This suggests that the patched transformer architectures use the information content in the data more efficiently.
@@
ViT NO $5.41\cdot 10^{-3}$ $1.97\cdot 10^{-2}$ $1.59\cdot 10^{-2}$
FANO $4.41\cdot 10^{-3}$ $1.68\cdot 10^{-2}$ $8.56\cdot 10^{-3}$
FNO $7.01\cdot 10^{-3}$ $3.31\cdot 10^{-2}$ $2.52\cdot 10^{-2}$
@@ 6.2.2 Darcy Flow
$C=48^{2}\bigl{(}-\Delta+12^{2}I\bigr{)}^{-2},$
@@
For the low resolution setting, we train the FNO and transformer neural operator architectures using $3600$ independent samples of resolution $64\times 64$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. On the other hand, the transformer neural operator is parametrized by $128$ channels in the encoder and is trained using a batch size of $2$ and learning rate of $10^{-3}$. On this problem we train both neural operators using the relative $H^{1}$ loss.
For the high resolution setting, we train the FNO, ViT neural operator and Fourier attention neural operator architectures using $3600$ independent samples of resolution $416\times 416$; we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is again parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (see Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of $1$ and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 11 displays the high resolution test results obtained by applying the Fourier attention neural operator.
@@
In the context of the piecewise constant experimental setting, we train the FNO, ViT neural operator and Fourier attention neural operator architectures using $3600$ independent samples of resolution $416\times 416$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is again parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (see Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of $1$ and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 12 displays the high resolution test results obtained by applying the Fourier attention neural operator.
@@
In Figures 13 and 14 we display the results of applying the neural operator architectures for the lognormal and piecewise input settings, respectively, to test samples of different resolutions than the one used for training. The results demonstrate the zero-shot generalization to different resolutions capability of the transformer neural operator architectures, which does not require retraining of the models. In the lognormal setting the FNO exhibits invariance to discretization that is more stable to changing resolution than are the ViT neural operator and the Fourier attention neural operator.

### v2

@@ 5 Universal Approximation by Transformer Neural Operators
The existence of $W_{1},R_{1},R_{2},P_{1},P_{2},b_{1},b_{R},b_{R}^{\prime},b_{P},b_{P}^{\prime}$ so that $\mathsf{G}(\cdot\,;\theta)$ satisfies (35) then follows from Lanthaler et al. (2025, Theorem 2.1), which also involves the application of the universality result for two-layer neural networks of Pinkus (1999, Theorem 4.1).
The analysis in Lanthaler et al. (2025) allows the derivation of an analogous universal approximation theorem for functions belonging to Sobolev spaces. This generalization is the content of the next theorem. The proof follows easily by applying the same argument as in the proof of Theorem 22 and the result of Lanthaler et al. (2025, Theorem 2.2).
@@ 6 Numerical Experiments
In this section we illustrate, through numerical experiments, the capability of the transformer neural operator architectures described in Section 4. Throughout, we consider the supervised learning problem described by the data model in (3) and take the viewpoint of surrogate modeling to construct approximations of the operators $\mathsf{G}^{\dagger}$. In Subsection 6.1 we consider problems given by dynamical systems and ordinary differential equations, namely operators acting on functions defined on a one-dimensional time domain. In this context, we explore the use of the transformer neural operator for operator learning problems given by the Lorenz ‘63 dynamical system and a controlled ODE. On the other hand, in Subsection 6.2 we consider problems given by partial differential equations equations, namely operators acting on functions defined on a two-dimensional spatial domain. In this context we explore the use of the transformer neural operator and of the more efficient ViT neural operator and Fourier attention neural operator architectures for operator learning problems given by the Darcy flow and Navier-Stokes equations. The code repository for the experiments in this paper may be found at https://github.com/EdoardoCalvello/TransformerNeuralOperators.
@@
${dz}$ $=\bigl{(}xy-\beta z\bigr{)}{dt},$
with parameters set to $\sigma=10$, $\rho=28$ and $\beta=8/3$. We consider learning the operator $\mathsf{G}^{\dagger}:C\bigl{(}[0,T];\mathbbm{R}^{d_{x}}\bigr{)}\times\mathbbm{R}^{d_{y}}\times\mathbbm{R}^{d_{z}}\to C\bigl{(}[0,T];\mathbbm{R}^{d_{y}}\bigr{)}\times C\bigl{(}[0,T];\mathbbm{R}^{d_{z}}\bigr{)}$ given by
@@
The experimental settings we consider for the continuous time operator learning problem may be viewed through the lens of data assimilation as smoothing problems (Sanz-Alonso et al., 2023). Indeed, the mappings to be learned represent the recovery of an unobserved trajectory $\{z(t)\}_{t\in{[0,T]}}$ conditioned on the observation of a whole trajectory $\{u(t)\}_{t\in{[0,T]}}$.
Figure 7: The panel displays the performance in relative $L^{2}$ errors (in log-scale) of the transformer neural operator when applied to the Lorenz ‘63 and controlled ODE operator learning test sets of different resolutions at inference time (without retraining). All models were parametrized by $d_{\textrm{model}}=128$ and 66 encoder layers, and were trained with $8000$ trajectories with $T=2$, sampled according to $\Delta t=0.01$.
@@
In other words, the second half of the testing domain is up-sampled by a factor of 2 compared to the training discretization. As seen qualitatively in Figure 8 testing on the non-uniform discretization based on a model trained on the uniform grid is successful. The use of different grids in test and train does lead to a larger error than the one incurred when they match (Figure 7), but the errors remain small. It is our continuum formulation of attention from Section B.1 that enables this direct generalization; indeed, as displayed in Figure 9, the transformer from Vaswani et al. (2017) when applied to the same setting yields larger errors than the transformer neural operator, due to the TNO’s reweighting based on the discretization of the grid.
@@
Figure 9: The panel displays the performance in relative $L^{2}$ errors of the transformer from Vaswani et al. (2017) when applied to the Lorenz ‘63 operator learning problem of recovering the unobserved $y$ trajectory. The results displayed concern a model trained on trajectories indexed on the index set (42) deployed on a test of trajectories indexed on (41). In each column, the top plot shows the input data and the bottom compares the prediction and ground truth; the left column shows the testing example achieving a median error, and the right columns shows the test example achieving the worst error.
@@
In this subsection we study two different 2D operator learning problems – flow in a porous medium, governed by the Darcy model, and incompressible viscous fluid mechanics governed by the Navier–Stokes equation, in particular, a Kolmogorov flow. We apply the three architectures developed in Sections 2 and 3, concentrating on the patched architectures. In the first Subsection 6.2.1 we report the results obtained by applying the neural operators described to an array of operator learning problems, where the domain of the functions is two-dimensional i.e. $D\subset\mathbbm{R}^{2}$. In Subsection 6.2.2 we describe the settings of the two Darcy flow operator learning problems considered, along with the parametrization and training details for the various architectures. Similarly in Subsection 6.2.3, we describe the Kolmogorov flow operator learning problem and relevant implementation details. Subsections 6.2.2 and 6.2.3 also contain further numerical results illustrating the behaviour of the proposed transformer neural operators, and discretization invariance in particular.
@@ 6.2.1 Results
The results presented concern the two variants of Darcy flow, with lognormal and piecewise constant inputs, and the Kolmogorov flow. The operator learning architectures we deploy include the three transformer-based methods introduced in this paper, and the implementations of the Fourier neural operator from the “NeuralOperator” library (Li et al., 2021; Kovachki et al., 2023), the Galerkin transformer from the “galerkin-transformer” library (Cao, 2021), and the AFNO (Guibas et al., 2022) from the “physics-nemo” library (PhysicsNeMo Contributors, 2023). We note that in order to obtain a numerical comparison of the attention mechanisms themselves, in the context of the Galerkin transformer architecture we only use the Galerkin transformer encoder, with lifting and projections achieved by linear layers as in TNO. This allows for a more direct comparison between the Fourier attention from Cao (2021) and our continuum attention, without potential expressivity gain from additional architectural components. In the following discussion, we do not include experiments performed using the Codomain-attention neural operator (Rahman et al., 2024) in the high-resolution settings: the memory usage with the implementation from the “NeuralOperator” library did not allow for comparably sized models, in terms of parameters; the performance of the models we trained was hence incomparably inferior to the others.
Our first results compares the transformer neural operator from Subsection 4.2 with FNO, AFNO and the Galerkin transformer (employing Fourier attention) on the lognormal Darcy Flow problem in a low $64\times 64$ resolution setting. Table 1 provides a brief summary of the test relative $L^{2}$ errors obtained. It is notable that the transformer architecture obtains the best performance with an order of magnitude fewer parameters. However, since higher resolution renders the architecture from Subsection 4.2 impractical in dimensions $d\geq 2$ the remainder of our experiments use the patched ViT and Fourier attention neural operators from Subsections 4.3 and 4.4 respectively.
@@
Transformer NO $5.98\cdot 10^{5}$ $1.19\cdot 10^{-2}$
FNO $5.70\cdot 10^{6}$ $1.96\cdot 10^{-2}$
AFNO $1.10\cdot 10^{6}$ $2.57\cdot 10^{-2}$
Galerkin Transformer $6.04\cdot 10^{5}$ $7.30\cdot 10^{-2}$
@@
We now consider training two patched architectures on the Darcy Flow problem with piecewise constant inputs and the Kolmogorov Flow problem; in both cases we use a $416\times 416$ resolution setting. We compare cost versus accuracy for the different neural operators implemented, defining cost in terms of both number of parameters and number of FLOPS. In Table 2 we present the FLOPS for each method; details of the derivation are provided in Appendix C. We note that the parameter scaling for the transformer neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\bigr{)}$, while the scaling for the FNO and the Fourier attention neural operator is $\mathcal{O}\bigl{(}d_{\textrm{model}}^{2}\cdot k_{\textrm{max}}\bigr{)}$, where $k_{\textrm{max}}$ is the total number of Fourier modes used. The parameter scaling for the ViT neural operator is given by $\mathcal{O}\bigl{(}d_{\textrm{model}}\cdot k_{\textrm{max}}+d_{\textrm{model}}^{2}\bigr{)}$. We note that the parameter scaling for the AFNO architecture is given by $\mathcal{O}\bigl{(}N^{2}d_{\textrm{model}}+d_{\textrm{model}}^{2}\bigr{)}$, where $b$ is a chosen integer hyperparameter, and where the $N^{2}d_{\textrm{model}}$ appears because of the learnable positional encoding. Further details are provided in Appendix C.
@@
AFNO $\mathcal{O}\bigl{(}d_{\textrm{model}}N\log(\sqrt{N})+Nd^{2}_{\textrm{model}}/b\bigr{)}$
@@
Table 2: Evaluation cost (measured in FLOPS) for the three proposed transformer neural operators compared to the Fourier neural operator (Li et al., 2021) and Adaptive Fourier neural operator (Guibas et al., 2022). We include the scaling of the AFNO as implemented for our use case, i.e. not employing patching and not employing mode truncation. We give further details on the derivation of these scalings in Appendix C.
Figures 10 and 11 display the cost-accuracy trade-off. Table 3 provides the test relative $L^{2}$ errors values, depicted in the cost-accuracy analysis figures, obtained with the models with the highest parameter count. Figure 10 demonstrates the parameter-efficiency of the ViT neural operator in comparison with the FNO, AFNO and Fourier attention neural operator. Once cost is defined in terms of FLOPS it is apparent that both patched transformer architectures outperform the FNO; furthermore in the Kolmogorov flow problem, the FNO exhibits accuracy-saturation, caused by the finite data set, whereas the patched architectures are less affected, in the ranges we test here. This suggests that the patched transformer architectures use the information content in the data more efficiently. The AFNO is comparable in terms of parameter count to FANO but more costly than ViTNO. On the other hand, even with no patching, the AFNO is cheaper in terms of evaluation complexity. Both ViTNO and FANO outperform the AFNO in terms of parameter-accuracy tradeoff in the context of the Darcy flow problem. For the same problem ViTNO and FANO bring an improvement in accuracy, albeit at a higher computational complexity. In the context of the Kolmogorov flow problem, ViTNO and FANO are competitive with AFNO in parameter-accuracy tradeoff. However for this problem, the AFNO can achieve higher accuracy with a lower complexity. It is important to note that because of the learnable positional embedding, and potential strided convolutions (if using patching), the AFNO is not a neural operator. This means that applying this architecture to different resolutions would require retraining, thus significantly increasing computational complexity. This is a fundamental advantage of the other architectures as their parameters may be deployed across different resolutions for training and inference. We recall the example of GenCast (Price et al., 2025), where a medium-range weather forecasting model is trained using a neural operator at low resolutions, and then finetuned at higher resolutions, thus drastically reducing the cost for such a large scale application.
@@
ViT NO $7.68\cdot 10^{-3}$ $9.72\cdot 10^{-3}$ $1.07\cdot 10^{-2}$
FANO $8.39\cdot 10^{-3}$ $1.38\cdot 10^{-2}$ $6.34\cdot 10^{-3}$
FNO $2.38\cdot 10^{-2}$ $3.78\cdot 10^{-2}$ $1.52\cdot 10^{-2}$
AFNO $2.66\cdot 10^{-2}$ $3.09\cdot 10^{-2}$ $3.99\cdot 10^{-3}$
@@ 6.2.2 Darcy Flow
$C=12^{2}\bigl{(}-\Delta+6^{2}I\bigr{)}^{-2},$
@@
For the low resolution setting, we train the FNO, AFNO, Galerkin transformer and TNO architectures using $3600$ independent samples of resolution $64\times 64$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. On the other hand, the transformer neural operator is parametrized by $128$ channels in the encoder and is trained using a batch size of 22 and learning rate of $10^{-3}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. The Galerkin transformer uses the Fourier type attention from Cao (2021), with quadratic complexity, making it prohibitive for higher resolutions. It uses $4$ layers of $128$ hidden channels and is trained using a learning rate of $10^{-3}$ and batch size of $2$. On this problem we train all neural operators using the relative $H^{1}$ loss.
For the high resolution setting, we train the AFNO, FNO, ViT neural operator and Fourier attention neural operator architectures using $3600$ independent samples of resolution $416\times 416$; we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is again parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (see Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of $1$ and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 12 displays the high resolution test results obtained by applying the Fourier attention neural operator.
@@
In the context of the piecewise constant experimental setting, we train the AFNO, FNO, ViT neural operator and Fourier attention neural operator architectures using $3600$ independent samples of resolution $416\times 416$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is again parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. Both the ViT neural operator and Fourier attention neural operator are parametrized by $128$ channels in the transformer encoder. The integral operator in the ViT NO is parametrized by $12$ Fourier modes in each dimension while the integral operators appearing in the Fourier attention neural operator are parametrized by $9$ Fourier modes in each dimension. Both architectures employ a final spectral convolution layer (see Remark 21) that is parametrized by $64$ Fourier modes in each dimension. Both of these neural operators are trained using a batch size of $1$ and learning rate of $10^{-3}$. On this problem we train all the neural operators with the relative $H^{1}$ loss. Figure 13 displays the high resolution test results obtained by applying the Fourier attention neural operator.
@@
In Figures 14 and 15 we display the results of applying the neural operator architectures for the lognormal and piecewise input settings, respectively, to test samples of different resolutions than the one used for training. The results demonstrate the zero-shot generalization to different resolutions capability of the transformer neural operator architectures, which does not require retraining of the models. In the lognormal setting the FNO exhibits invariance to discretization that is more stable to changing resolution than are the ViT neural operator and the Fourier attention neural operator.